If (3,N) is a soluton of Ax+3=6 and (n,5) is a solution of 5x+y=20, what is the value for A

Answer by stanbon(75887)

If (3,N) is a soluton of Ax+3=6 and (n,5) is a solution of 5x+y=20,
what is the value for A ?
-----
A*3 + 3 = 6
3A = 3
A = 1
-----------------
5n+5 = 20
5n = 15
n = 3
---------------
Cheers,
Stan H.
--------------